# Good News for Polynomial Root-finding

### v35

@@ Abstract
Univariate polynomial root-finding has been studied for four millennia and very intensively in the last decades. Our novel near optimal root-finders approximate all zeros of a polynomial $p$ almost as fast as one accesses its coefficients with the precision required for the solution within a prescribed error bound.  [^1] Under Las Vegas randomization we allow output errors but detect them at a dominated computational cost. Furthermore, our root-finders can be applied to a black box polynomial, defined by an oracle (that is, black box subroutine) for its evaluation rather than by its coefficients. [^2] The pioneering root-finder by Louis and Vempala at FOCS 2016 is the only other known black box root-finder, but it only approximates the absolutely largest root of a real-rooted polynomial. Such root-finders are particularly fast for polynomials that can be evaluated fast such as the sum of a few shifted monomials of the form $(x-c)^{d}$. Our divide and conquer algorithm of ACM STOC 1995 is the only other known near optimal polynomial root-finder, but it extensively uses the coefficients, is quite involved, and has never been implemented, while according to extensive numerical experiments with standard test polynomials, already initial implementations of our new root-finders compete with user’s choice package of root-finding subroutine MPSolve and supersede it more and more significantly where the degree of a polynomial grows large. [^3] The root-finder of 1995 is deterministic, while our current ones apply Las Vegas randomization, that is, allow output errors but detect them at a dominated computational cost. Randomization enables us to avoid (with a high probability) division by near vanishing values in our computations with bounded precision. Our root-finders are readily extended to support approximation of the eigenvalues of a matrix within a record Las Vegas expected bit operation time bound. Our auxiliary algorithms and techniques for computations with black box polynomials can be of independent interest.
@@
polynomial computations, matrix eigenvalues, complexity, polynomial zeros, black box polynomials, root-squaring, symbolic-numeric computing.
@@
1 Introduction
@@ 1.1 Polynomial Root-finding: Problems, Impact, and Earlier Studies
According to [7, 95, 96, 80], univariate polynomial root-finding, that is, the solution of a polynomial equation $p(x)=0$, had been the central problem of Mathematics and Computational Mathematics for about four millennia since Sumerian times; it was still a central problem well into the 19th century. In its long way from its rudimentary level to its modern version, this study has been responsible for the introduction of mathematical rigor and notation as well as the concepts of irrational, negative, and complex numbers, algebraic groups, rings, and fields etc. Polynomial root-finding is still an important subject of modern computations, in particular for computer algebra, signal processing, geometric modeling, control, and financial mathematics [77, 79, 108].
@@
Already about 500 BC it was proved in Pythagorean school that the roots of the equation $x^{2}=2$ were irrational radicals, while even with the help of radicals one cannot express the roots rationally if the degree $d$ of $p(x)$ exceeds 4, according to Ruffini (1799,1813) and Abel (1824); moreover, one cannot do this already for specific equations such as $x^{5}-x-1=0$ or $x^{5}-4x-2$, according to Galois theory of 1830, 1832. Weyl in his constructive proof of Fundamental Theorem of Algebra in 1924 in [144], however, rationally approximated all $d$ zeros of $p$ within any fixed positive error tolerance. This progress, the advance of computer technology, and various applications to modern computations have led to the study of the following
Root-finding problems: Given the $d+1$ complex coefficients of a polynomial $p(x)$ of a degree $d$, approximate within a fixed error tolerance $\epsilon$ either (i) its every zero (aka every root of the equation $p(x)=0$) or (ii) its every zero lying in a fixed Region of Interest $\mathbb{S}$ [^4] We count the zeros of $p$ with multiplicities and do not distinguish multiple roots from root clusters.
Hundreds of efficient root-finders have been proposed [77, 79] and keep appearing (see [118, 21, 38, 67, 73, 123, 104, 22, 23, 61, 106, 107, 72, 57, 59, 78, 60, 56, 111], and the bibliography therein). Smale in [129] proposed to study Problem 1 in terms of arithmetic complexity and computational precision and later (jointly with Blum, Cucker, and Shub) (cf. [13, 28]) linked this study to rounding errors and conditioning of the roots. In [130] Schönhage proposed to estimate the complexity of polynomial root-finders in terms of bit operations involved and presented an algorithm that approximated all zeros of $p$ within $\epsilon=1/2^{b}$ by using $\tilde{O}(bd^{3})$ bit operations, that is, $bd^{3}$ bit operations up to a polylogarithmic factor in $bd$. Intensive study in the 1980s and 1990s has culminated in [92] (cf. also [93, 99] and surveys in [95, 96]) with our divide and conquer root-finder that decreased Schönhage’s record to $\tilde{O}((b+d)d^{2})$. This near optimal bound, that is, optimal up to polylogarithmic factors, implies that the root-finder runs nearly as fast as one accesses the coefficients with the precision required for the task.
If we only consider polynomials whose zeros and tiny clusters of zeros are pairwise sufficiently isolated (as we specify in Def. B.1 of Sec. B), then the known lower bounds on the computational precision and the bit operation complexity of root-finding decrease by a factor of $d$, and so do the upper bounds supported by the algorithms of [92, 98] as well; hence the algorithms remain near optimal.
1.2 Reduction of root-finding problems to the case of the unit disc
Hereafter ${}^{T}$ shows transposition of a vector or a matrix, $\mathbb{R}^{k}$ and $\mathbb{C}^{k}$ denote the spaces of the column vectors $(v_{1},\dots,v_{k})^{T}$ of length $k$ with real and complex coordinates, respectively, and $D(c,\rho):=\{x:~{}|x-c|\leq\rho\}$ and $C(c,\rho):=\{x:~{}|x-c|=\rho\}$ denote the disc of radius $\rho>0$ centered at $c\in\mathbb{C}$ and its boundary circle, respectively.
Next we link the above Root-finding problem (i) to its case (ii) where the Region of Interest $\mathbb{S}$ is the unit disc $D(0,1)$.
Problem 0: Root approximation in the unit disc.
INPUT: positive $\epsilon\in\mathbb{R}$, $d\in\mathbb{N}$, and $(p_{0},p_{1},\dots,p_{d})^{T}\in\mathbb{C}^{d}$, defining the polynomial
@@ 1.2 Reduction of root-finding problems to the case of the unit disc
APPROXIMATE within $\epsilon$ all its zeros lying in the unit disc $D(0,1)$. [^5] We will also refer to $z_{1},\dots,z_{d}$ as “roots of equation $p=0$” or just “roots”.
(i) Two ways from the zeros of $p$ lying in the unit disc to all its zeros.
@@
(a) Given an algorithm $\mathcal{A}$ that solves Problem 0, we can compute approximations $\tilde{y}_{j}$ within $\epsilon$ to the reciprocals $y_{j}=1/x_{j}$ of all zeros $x_{j}$ of $p$ lying outside the disc $D(0,1)$ [130] – simply by applying this algorithm to the reverse polynomial
@@
Write $\Delta_{j}:=\tilde{y}_{j}-y_{j}$ and deduce that $\frac{1}{y_{j}}-\frac{1}{\tilde{y}_{j}}=\frac{\Delta_{j}}{\tilde{y}_{j}~{}y_{j}}$, and so $\frac{\tilde{y}_{j}}{y_{j}}-1=\frac{\Delta_{j}}{y_{j}}=\Delta_{j}x_{j}$. Recall that $|\Delta_{j}|\leq\epsilon$ by assumption and obtain that $|\frac{\tilde{y}_{j}}{y_{j}}-1|\leq\epsilon|x_{j}|$, thus arriving at the following
@@
(b) If in addition to an algorithm $\mathcal{A}$ we are given a value
$R\geq\max_{j=1}^{d}\{|x_{j}|\},$ (1.3)
then we can scale the variable $x\mapsto y=\frac{x}{R}$, $p(x)\mapsto p(y)=p(\frac{x}{R})$ and then approximate every root $x_{y}$, for $j=1,\dots,d$, within $\epsilon R$ by solving Problem 0.
(ii) From the zeros of $p$ in the unit disc to all zeros in a Region of Interest.
For any pair of complex $c$ and positive $\rho$ we can reduce root-finding in the disc $D(c,\rho)$ to Problem 0 by means of applying the linear map
which shifts the variable $x$ by $c$ and then scales $x-c$ by $1/\rho$.
Now, given any bounded region of interest $\mathbb{S}$ on the complex plane, we can successively (a) cover it with a minimal disc $D(c,\rho)$ (see Remark 1.4 of Sec. 1.6.2 for closer covers), (b) map this disc into the disc $D(0,1)$, (c) solve Problem 0 for the resulting polynomial $t(x)$ of (1.4), and finally, (d) apply map (1.4) in the reverse direction to obtain approximations to all zeros of $p$ lying in the intersection of the $\epsilon$-neighborhood $\mathbb{S}_{\epsilon}$ of the region $\mathbb{S}$ with the covering disc $D(c,\rho)$.
1.3 Black Box Polynomial Root-finding and Eigenvalue Approximation: Problem Statements
A nontrivial variant of Newton’s iterations in the FOCS 2016 paper [73] by Louis and Vempala approximates within a fixed positive $\epsilon\leq 1/2$ an absolutely largest zero $z$ of a real-rooted polynomial $p$ of a degree $d$ at the cost of its evaluation with a precision in $O((b+\log(R))\log(R~{}d)))$ at $O((b+\log(R))\log(d))$ points for $b:=\log_{2}(R/\epsilon)$ and a fixed $R$ satisfying (1.3). While the previous polynomial root-finders only worked for polynomials given in the monomial basis, that is, with their coefficients, the algorithm of [73] never involves them and works for a black box polynomial $p$ – given with an oracle (black box subroutine) for its evaluation. Next we modify Problem 0 accordingly.
Problem 1: Black Box Polynomial Root Approximation.
INPUT: a positive $\epsilon\in\mathbb{R}$, $d\in\mathbb{N}$, and a $d$th degree black box polynomial $p$.
APPROXIMATE within $\epsilon$ the zeros of $p$ lying in the unit disc $D(0,1)$.
By applying their root-finder to the characteristic polynomial $p(x)=\det(xI-M)$ of a $d\times d$ symmetric matrix $M$, Louis and Vempala approximated its absolutely largest eigenvalue by using a record Las Vegas expected number of bit operations. [^6] Under Las Vegas randomization one allows output errors but detects them at a dominated computational cost. Accordingly, we extend Problem 0 as follows:
@@ 1.3 Black Box Polynomial Root-finding and Eigenvalue Approximation: Problem Statements
INPUT: positive $\epsilon\in\mathbb{R}$, $d\in\mathbb{N}$, and a $d\times d$ matrix $M$.
APPROXIMATE within $\epsilon$ the eigenvalues of $M$ lying in the unit disc $D(0,1)$.
The algorithms for Problems 0, 1, and 2 may output multiple approximations within $\epsilon$ to the same roots and eigenvalues, but in the next modification we avoid such a confusion.
Problems $0^{*}$, $1^{*}$, $2^{*}$: Isolation of Clusters of the Zeros of $p$ and the Eigenvalues of $M$. Under the assumptions of Problems 0, 1, and 2 compute at most $d$ disjoint discs of radii at most $\epsilon$ (hereafter $\epsilon$-discs) covering all zeros or eigenvalues that lie in the disc $D(0,1)$ and also compute the number of the zeros or eigenvalues lying in each $\epsilon$-disc.
@@
If we collapse all zeros of $p$ lying in every $\epsilon$-disc into a single multiple zero placed in the center of the disc, then these zeros will lie at the distances greater than $2\epsilon$ from each other. If polynomial $p$ is square-free (having only simple zeros) and has integer coefficients, then we would arrive at the solution of the well-known Problem of Root-isolation (see Problem ISOL and Remark C.4 in Sec. C.5). These comments can be immediately extended to Problem $*{}^{*}$.
1.4 Some Benefits of Black Box Polynomial Root-finders

### v36

# Fast Polynomial Root-Finding and Approximation of the Eigenvalues of a Matrix

@@ Abstract
Univariate polynomial root-finding has been studied for four millennia and very intensively in the last decades. Our new near-optimal root-finders approximate all zeros of a polynomial $p$ almost as fast as one accesses its coefficients with the precision required for the solution within a prescribed error bound. Furthermore, our root-finders can be applied to a black box polynomial, defined by an oracle (black box subroutine) for its evaluation rather than by its coefficients. Due to this feature our root-finders (i) support approximation of the eigenvalues of a matrix in a record Las Vegas expected bit operation time and (ii) are particularly fast for a polynomial that can be evaluated fast such as the sum of a few shifted monomials of the form $(x-c)^{d}$ or a Mandelbrot-like polynomial defined by a recurrence. Our divide and conquer algorithm of ACM STOC 1995 is the only other known near-optimal polynomial root-finder, but it extensively uses the coefficients, is quite involved, and has never been implemented, while according to extensive numerical experiments with standard test polynomials, already a slower initial implementation of our new root-finders competes with user’s choice package of root-finding subroutines MPSolve and supersedes it more and more significantly as the degree of a polynomial grows large. Our auxiliary algorithms and techniques for computations with black box polynomials can be of independent interest.
@@
polynomial root-finding, matrix eigenvalues, computational complexity, black box polynomials, symbolic-numeric computing.
@@
PART I: BACKGROUND AND OUR PROGRESS
1 State of the Art and our Main Results
In this part of the paper we state our main results and outline our algorithms. We present them in some detail in Parts II, III, and IV, which can be read independently of each other.
@@ 1.1 Polynomial Root-finding: Problems, Impact, and Earlier Studies
According to [7, 99, 100, 83], univariate polynomial root-finding, that is, the solution of a polynomial equation $p(x)=0$, had been the central problem of Mathematics and Computational Mathematics for about four millennia since Sumerian times; it was still a central problem well into the 19th century. In a long way from its rudimentary level to modern version, its study has been responsible for the introduction of mathematical rigor and notation as well as the concepts of irrational, negative, and complex numbers, algebraic groups, rings, and fields etc.
@@
Already about 500 BC it was proved in Pythagorean school that the roots of the equation $x^{2}=2$ were irrational radicals, while even with the help of radicals one cannot express the roots rationally if the degree $d$ of $p(x)$ exceeds 4, according to Ruffini (1799,1813) and Abel (1824); moreover, one cannot do this already for specific equations such as $x^{5}-x-1=0$ or $x^{5}-4x-2$, according to Galois theory of 1830, 1832.
Nevertheless, in 1924 in [149], Herman Weyl published his constructive proof of the Fundamental Theorem of Algebra where he rationally approximated within any fixed positive error tolerance all roots of $p(x)=0$ lying in a fixed square in the complex plane, that is, all roots if the square is large enough. This was the solution of the following Problem 0 of modern polynomial root-finding:
INPUT: $d+1$ complex coefficients $p_{0},p_{1},\dots,p_{d}$ of a polynomial
@@
a positive error tolerance $\epsilon=1/2^{b}>0$, an integer $m$, $1\leq m\leq d$, and a bounded domain $\mathbb{D}$ in the complex plane, e.g., a disc or a rectangle, such that $z_{j}\in\mathbb{D}$ if and only if $j\leq m$.
OUTPUT: approximations $\tilde{z}_{j}\approx z_{j}$, $j=1,\dots,m$, to the $m$ zeros of $p(x)$ (aka roots of $p(x)=0$) [^1] Hereafter we freely say just “root” or “zero” for short, always count them with their multiplicities, and do not distinguish multiple roots from pairwise isolated root clusters of diameters less than $\epsilon$. lying in the domain $\mathbb{D}$ such that
$|\tilde{z}_{j}-z_{j}|\leq\epsilon~{}{\rm for}~{}j=1,\dots,m.$ (1.2)
Example 1.1.
[Approximation of all roots.] Fix $R\geq\max_{j}|z_{j}|$ (see [52] for simple expressions for such $R$) and let $\mathbb{D}$ denote the disc $D(0,R):=\{x:~{}|x\leq R|\}$, centered in the origin and having radius $R$. Then Problem 0 requires us to approximate all $d$ roots.
Example 1.2.
[Numerical approximation of real roots.] Root-finding in the real line or its segment (real root-finding) is highly important because in many applications, e.g., to optimization in algebraic geometry and geometric modeling, only real roots of a polynomial are of interest, and typically they are much less numerous than all $d$ complex roots (cf. Kac [66], Edelman and Kostlan [39]). More generally consider Problem 0 where $\mathbb{D}$ denotes the minimal rectangle in the complex plane that covers the $\epsilon/\sqrt{2}$-neighborhood of a fixed real line segment. To solve this problem we should approximate both real and near-real roots, that is, their perturbations due to small input and rounding errors. The known real root-finders (cf. [80, Chs. 2 and 3]) do not address this important issue.
Major motivation for modern study of Problem 0 comes from application to approximation of the eigenvalues of a matrix (see our Secs. 1.3 and 1.5) and the solution of a multivariate polynomial system of equations [29], both problems being highly important for the theory and practice of modern computations. [82, Introduction] specifies some applications of polynomial root-finding to computer algebra, signal processing, geometric modeling, control, and financial mathematics. Furthermore, Problem 0 for a domain $\mathbb{D}$ containing all roots is closely linked to polynomial factorization (see Secs. 24 and 25), which has applications to time series analysis, Wiener filtering, noise variance estimation, co-variance matrix computation, and the study of multichannel systems [150, 18, 8, 35, 36, 147].
Hundreds of efficient root-finders have been proposed [80, 82] and keep appearing (see [122, 22, 40, 70, 76, 127, 108, 23, 24, 64, 110, 111, 75, 60, 62, 81, 63, 59, 115], and the bibliography therein). In [133] Smale proposed to study Problem 0 in terms of arithmetic complexity and computational precision and later, jointly with Blum, Cucker, and Shub (cf. [14, 30]), linked this study to rounding errors and conditioning of the roots. In [134] Schönhage proposed to estimate the complexity of polynomial root-finders in terms of the number of bit operations involved and presented an algorithm that solved Problem 0 for a domain that contains all roots by using $\tilde{O}(bd^{3})$ bit operations, [^2] Here and hereafter we assume that $b\geq d$, recalling that error amplification factors $2^{cd}$ are inevitable in polynomial root-finding, e.g., in routine shift of the variable. that is, $bd^{3}$ bit operations up to a polylogarithmic factor in $bd$.
His work prompted intensive study of root-finding by researchers in the Theory of Computing in the 1980s and 1990s, culminated at ACM STOC 1995 with our near-optimal divide and conquer root-finder of [96] (also see its journal versions [97, 103] and surveys [99, 100]); it supports bound $\tilde{O}(bd^{2})$ and showed the information lower bound $0.25(d+1)bd$ (see Cor. 24.2 in Sec. 24.2).
If we only consider polynomials whose zeros and tiny clusters of zeros are pairwise sufficiently isolated, as we specify in Def. 24.1 of Sec. 24, then the lower bounds on the computational precision and the bit operation complexity of root-finding decrease by a factor of $d$ (see Sec. 24), and so do the upper bounds supported by the algorithm of [96, 97, 103] as well; hence the algorithm remains near-optimal (see Thms. 24.1 and 24.4 and Cor. 24.3 in Sec. 24).
1.2 Reduction to the case of the unit disc
Hereafter T shows transposition of a vector or a matrix, $\mathbb{R}^{k}$ and $\mathbb{C}^{k}$ denote the spaces of the column vectors $(v_{1},\dots,v_{k})^{T}$ of length $k$ with real and complex coordinates, respectively, and $D(c,\rho):=\{x:~{}|x-c|\leq\rho\}$ and $C(c,\rho):=\{x:~{}|x-c|=\rho\}$ denote the disc of radius $\rho>0$ centered at $c\in\mathbb{C}$ and its boundary circle, respectively. Furthermore, let Problem $\widehat{0}$ denote Problem 0 where the domain $\mathbb{D}$ is the unit disc $D(0,1)$; we can extend its solution to root-finding in the disc $D(c,\rho)$ for any pair of complex $c$ and positive $\rho$ by means of applying the linear map
which shifts the variable $x$ by $c$ and then scales $x-c$ by $1/\rho$.
Next we reduce Problem 0 for any domain $\mathbb{D}$ to Problem $\widehat{0}$.
(i) Two ways from the roots in the unit disc to all roots. Given a value
$R\geq\max_{j=1}^{d}\{|x_{j}|\}$ (1.4)
(see [52] for simple expressions for such $R$), we can scale the variable $x\mapsto y=\frac{x}{R}$, $p(x)\mapsto p(y)=p(\frac{x}{R})$ and approximate every root $x_{y}$, for $j=1,\dots,d$, within $\epsilon R$ by solving Problem $\widehat{0}$ for $p(y)$.
@@ 1.2 Reduction to the case of the unit disc
Schönhage in [134] reduced approximation of all $d$ roots to Problem $\widehat{0}$ without assuming (1.4) or using (1.3) as follows. Given an algorithm $\mathcal{A}$ that solves Problem $\widehat{0}$, compute also approximations $\tilde{y}_{j}$ within $\epsilon$ to the reciprocals $y_{j}=1/z_{j}$ of all roots lying outside the disc $D(0,1)$ – simply by applying this algorithm to the reverse polynomial
@@
Write $\Delta_{j}:=\tilde{y}_{j}-y_{j}$ and deduce that $\frac{1}{y_{j}}-\frac{1}{\tilde{y}_{j}}=\frac{\Delta_{j}}{\tilde{y}_{j}~{}y_{j}}$, and so $\frac{\tilde{y}_{j}}{y_{j}}-1=\frac{\Delta_{j}}{y_{j}}=\Delta_{j}z_{j}$. Recall that $|\Delta_{j}|\leq\epsilon$ by assumption and obtain that $|\frac{\tilde{y}_{j}}{y_{j}}-1|\leq\epsilon|z_{j}|$, thus arriving at the following
@@
(ii) From roots in the unit disc to roots in a bounded domain. Given a bounded domain $\mathbb{D}$ on the complex plane, we can successively (a) cover it with a minimal disc $D(c,\rho)$, (b) map this disc into the disc $D(0,1)$, (c) solve Problem $\widehat{0}$ for the resulting polynomial $t(x)$ of (1.3), and finally, (d) use the map (1.3) again – this time to approximate all roots lying in the intersection of the disc $D(c,\rho)$ with the domain $\mathbb{D}$. In this way we also approximate extraneous roots lying in $D(c,\rho)$ but not in $\mathbb{D}$; in Sec. 2.2 we control the computational cost of that additional work.
1.3 Black box root-finders; the isolation of roots, eigenvalues, and their clusters
A nontrivial variant of Newton’s iterations by Louis and Vempala [76], at FOCS 2016, approximates within a fixed positive $\epsilon\leq 1/2$ an absolutely largest zero of a real-rooted polynomial $p$ of a degree $d$ at the cost of its evaluation with a precision in $O((b+\log(R))\log(R~{}d)))$ at $O((b+\log(R))\log(d))$ points for $b:=\log_{2}(R/\epsilon)$ and a fixed $R$ satisfying (1.4). While the previous polynomial root-finders have been devised for polynomials given in the monomial basis, that is, with their coefficients, the algorithm of [76] never involves them and works for a black box polynomial $p$ – given with an oracle (black box subroutine) for its evaluation. [76] shows benefits of extension of such algorithms to approximation of the eigenvalues of a matrix. Next we state the relevant tasks.
Problem 1 (Black Box Polynomial Root Approximation) is Problem 0 for a black box polynomial $p$.
@@ 1.3 Black box root-finders; the isolation of roots, eigenvalues, and their clusters
INPUT: $\epsilon=1/2^{b}>0$, integers $m$ and $d$ such that $0<m\leq d$, and a $d\times d$ matrix $M$, and a complex domain $\mathbb{D}$ that contains precisely $m$ eigenvalues of $M$ counted with their multiplicities.
OUTPUT: Approximations within $\epsilon$ to each of all $m$ eigenvalues of $M$ lying in the domain $\mathbb{D}$.
The output of Problems 0, 1, and 2 may include multiple approximations within $\epsilon$ to the same root or eigenvalue, but the next modifications exclude such a confusion.
Problems $0^{*}$, $1^{*}$, $2^{*}$: Isolation of Clusters of the Zeros of a polynomial and the Eigenvalues of a matrix. Under the assumptions of Problems 0, 1, and 2 compute at most $d$ disjoint discs of radii at most $\epsilon$ (hereafter $\epsilon$-discs) covering all roots or eigenvalues that lie in the domain $\mathbb{D}$; also compute the number of the roots or eigenvalues lying in each $\epsilon$-disc.
@@
If we collapse all zeros of $p$ lying in every $\epsilon$-disc into a single multiple zero placed in the center of the disc, then these zeros will lie at the distances greater than $2\epsilon$ from each other. If a polynomial $p$ is square-free (having only simple zeros) and has integer coefficients, then Problem 0 turns into the well-known Problem of Root-isolation (see Problem ISOL and Remark 25.4 in Sec. 25.5). These comments can be immediately extended to Problems 1 and 2.
1.4 More on Black Box Polynomial Root-finders
1.4.1 Why we should use them
One can perform black box evaluation at $q>d$ points $x_{1},\dots,x_{q}$ and then interpolate to $p$. The bit operation cost of interpolation, however, is unbounded unless the $q$ points are well isolated from the roots, but it is a challenge to yield such isolation. Furthermore, even if the isolation is ensured, the cost of interpolation exceeds the cost of the solution of Problems 0, 0, 1 ,1, 2, and 2 in the cases where domain $\mathbb{D}$ contains only a small number of roots.
Moreover, here are three examples of important benefits of using black box algorithms even where an input polynomial is given with its coefficients.
(i) Application of a black box root-finder to the characteristic polynomial $p(x)=\det(xI-M)$ of a $d\times d$ matrix $M$ only involves the values of this polynomial but not its coefficients, whose overall size exceeds the size (a norm) of the matrix $M$ by a factor of $d$. Consequently, extension of such a root-finder to Problem 2 saves a factor of $d$ in bit precision and bit operation count versus the known solutions of Problems 2 and 2, based on [134, Sec. 21].
(ii) Map (1.3) is a valuable root-finding tool, but its application is limited because it destroys sparsity and blows up computational precision and bit operation cost. Similar comments can be applied to maps (3.6) and (3.7) of Sec. 3.4. Black box polynomial root-finders keep all benefits of using those maps but avoid the latter notorious problems by involving no coefficients of $t(x)$ and $p_{h}(x)$ of these maps.